# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ 1.3 Forests of decision trees
Then $F$ is a decision forest over $V$ iff
@@
    for every $T\in\mathcal{F}$, $(T,\supseteq)$ defines a decision tree over $V_{T}$.
To verify the three propositions 1.4, 1.5, and 1.6, on can cite the results in [3] in combination with remark 1.2. In view of theorem 1.7, it suffices, however, to cite the results for the case of rooted trees only. This will be of interest at a later stage when analysing outcome existence and uniqueness for stochastic extensive forms.
@@
Theorem 1.7 offers a way of constructing decision forests from collections of decision trees. Let $\mathcal{F}_{0}$ be a non-empty set of decision trees over sets. For any $T_{0}\in\mathcal{F}_{0}$, let $V_{T_{0}}$ be its root. Then, let
@@
$F=\{x\times\{T_{0}\}\mid T_{0}\in\mathcal{F}_{0},x\in T_{0}\}.$
Then, $F$ is a decision forest over $V$. The set of connected components is given by
=$\mathcal{F}=\{\{x\times\{T_{0}\}\mid x\in T_{0}\}\mid T_{0}\in\mathcal{F}_{0}\}.$
Moreover, theorem 1.7 states that all decision forests over sets can be represented in this form.
For examples of decision forests over sets, the reader is referred to the following section, where decision forests whose connected components are indexed over the set of exogenous scenarios are considered. [^3] Also note the examples of decision trees over sets in [3] and the more extensive version in [5] in combination with the preceding remark 1.8.
@@ 2.1 Exogenous scenario spaces
For the remainder of the paper, an exogenous scenario space is a measurable space $(\Omega,\mathscr{A})$ such that $\Omega\neq\emptyset$. [^4] For more information on measure and probability theory, the reader is referred to the introductory texts [11, 28]. $\Omega$ describes the set of possible exogenous scenarios which replace the outcomes generated by the behaviour of a “nature” agent. $\mathscr{A}$ is the set of events alias collections of outcomes that can be measured by all relevant decision makers, in the sense, that their potential beliefs about events can be modelled via probability measures on $(\Omega,\mathscr{A})$. We formally follow the classical theory of subjective probability and its interpretation as a representation device of preferences in terms of expected utility, which goes back to Ramsey, de Finetti, Savage, among others. Accordingly, we think of an agent as an entity that has certain preferences over the result of the strategic interaction; according to the paradigm of the mentioned theory, this translates into a utility function over results and one or multiple probability measures alias beliefs on the exogenous scenario space. [^5] See [22, chapter 2] for a rigorous mathematical textbook presentation of the theory with multiple priors, going back to [23], alias robust preferences, underpinning robust control and risk measures.
@@
Why does one suppose $\mathscr{A}$ to be a $\sigma$-algebra? As an algebra of sets, $\mathscr{A}$ can be seen as a Boolean algebra and can therefore model basic logical operations on events. Moreover, in the probabilistic approach, we are only interested in those functions ${\mathbb{P}}\colon\mathscr{A}\to[0,1]$ compatible with the algebra structure [^6] The meaning of the phrase “compatible with the algebra structure” can be reduced to the statement: a) ${\mathbb{P}}(\Omega)=1$, and b) for all disjoint $A_{1},A_{2}\in\mathscr{A}$ we have ${\mathbb{P}}(A_{1}\cup A_{2})={\mathbb{P}}(A_{1})+{\mathbb{P}}(A_{2})$. that are $\sigma$-additive, which can be thought of as a continuity property, and it is a fundamental measure-theoretic result, that, for such ${\mathbb{P}}$, without loss of generality, $\mathscr{A}$ can be supposed to be closed under countable unions, that is, to be a $\sigma$-algebra (see [11, section 1.5]).
@@
If $\mathscr{A}$ can be generated by a countable partition $\mathscr{P}$ of $\Omega$, any probability measure on it is entirely described through the probabilities on these partitioning events. So, though all elements of $\mathscr{A}$ including unions of partition members are events, one can formally describe the situation by the set $\mathscr{P}$ and a countable family of positive numbers $(p_{A})_{A\in\mathscr{P}}$ adding up to $1$ and indexed by that set. However, in general, $\mathscr{A}$ need not be generated by a partition of $\Omega$. For instance, it may be that all singletons are events and have probability zero, as in the case of the Lebesgue measure on the unit interval $[0,1]$. Hence, from a general perspective, scenarios may be neither a relevant nor a sufficient description of exogenous data for decision makers, but it is the ($\sigma$-)algebra $\mathscr{A}$ of events that these agents are concerned about. Moreover, as famously illustrated by the Lebesgue measure, in case $\Omega$ is uncountable, there is a non-trivial trade-off between the fineness [^7] Recall that a $\sigma$-algebra $\mathscr{A}^{\prime}$ on $\Omega$ is finer than $\mathscr{A}$ iff $\mathscr{A}\subseteq\mathscr{A}^{\prime}$. of $\mathscr{A}$ and the amount of probability measures on it (see [11, section 1.7]). Hence, to describe the agent’s relation to exogenous data and admissible beliefs about their probability, it is crucial to specify $\mathscr{A}$. This is crucial, not the least because many relevant applications require an uncountable $\Omega$ (with a $\sigma$-algebra $\mathscr{A}$ that cannot be generated by a countable partition), be it the uniform distribution on $[0,1]$, be it continuous martingales omnipresent in finance (see, e.g., [17, 9]).
@@ 2.2 Stochastic decision forests
In light of the preceding argument, we model uncertainty about exogenous events by an exogenous scenario space $(\Omega,\mathscr{A})$ and suppose that some $\omega\in\Omega$ is realised, determining the relevant decision tree without being directly communicated to the decision makers. So, these agents do not necessarily know about the exogenous information telling in which tree they are while making choices based on the endogenous information given by moves, or “information sets” of moves – hence, there must be a structure of similarity among trees that can serve as a consistent basis for both endogenous and exogenous information revelation. This intuition is the basis of the following definition which is further interpreted and analysed in the sequel.
@@
A stochastic decision forest, in short sdf, over an exogenous scenario space $(\Omega,\mathscr{A})$ is a triple $(F,\pi,{\bm{\mathsf{X}}})$ consisting of:
@@
    a decision forest $F$ over some set $W$;
@@
      any element ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ is a section of moves defined on some event $D_{\bm{\mathsf{x}}}\in\mathscr{A}$, that is, it is a map ${\bm{\mathsf{x}}}\colon D_{\bm{\mathsf{x}}}\to X$ satisfying $\pi\circ{\bm{\mathsf{x}}}=\operatorname{id}_{D_{\bm{\mathsf{x}}}}$;
    2. (b)
      ${\bm{\mathsf{X}}}$ induces a covering of $X$, that is, $\{{\bm{\mathsf{x}}}(\omega)\mid{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}},\leavevmode\nobreak\ \omega\in D_{\bm{\mathsf{x}}}\}=X$;
    3. (c)
      if we consider the partial order $\geq_{\bm{\mathsf{X}}}$ on ${\bm{\mathsf{X}}}$ given by
      ${\bm{\mathsf{x}}}_{1}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{2}\quad\Longleftrightarrow\quad\Big{[}D_{{\bm{\mathsf{x}}}_{1}}\supseteq D_{{\bm{\mathsf{x}}}_{2}}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \forall\omega\in D_{{\bm{\mathsf{x}}}_{2}}\colon{\bm{\mathsf{x}}}_{1}(\omega)\supseteq{\bm{\mathsf{x}}}_{2}(\omega)\Big{]},$
      then for all ${\bm{\mathsf{x}}}_{1},{\bm{\mathsf{x}}}_{2}\in{\bm{\mathsf{X}}}$:
      $\exists\omega\in D_{{\bm{\mathsf{x}}}_{1}}\cap D_{{\bm{\mathsf{x}}}_{2}}\colon\leavevmode\nobreak\ {\bm{\mathsf{x}}}_{1}(\omega)\supseteq{\bm{\mathsf{x}}}_{2}(\omega)\qquad\Longrightarrow\qquad{\bm{\mathsf{x}}}_{1}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{2};$
    4. (d)
      for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, ${\bm{\mathsf{x}}}(\omega)$ is a root for all or no $\omega\in D_{\bm{\mathsf{x}}}$;
    5. (e)
      if $\bar{\bm{\mathsf{X}}}$ is a set sharing the previous four properties of ${\bm{\mathsf{X}}}$ that is refined by ${\bm{\mathsf{X}}}$ in that for all $\bar{\bm{\mathsf{x}}}\in\bar{\bm{\mathsf{X}}}$ there is $P_{\bar{\bm{\mathsf{x}}}}\subseteq{\bm{\mathsf{X}}}$ with $\bar{\bm{\mathsf{x}}}=\bigcup P_{\bar{\bm{\mathsf{x}}}}$, [^8] According to set-theoretic conventions, $\bar{\bm{\mathsf{x}}}=\bigcup P_{\bar{\bm{\mathsf{x}}}}$ means: $\bar{\bm{\mathsf{x}}}$ is a map with domain $\bigcup_{{\bm{\mathsf{x}}}\in P_{\bar{\bm{\mathsf{x}}}}}D_{\bm{\mathsf{x}}}$ and for all ${\bm{\mathsf{x}}}\in P_{\bar{\bm{\mathsf{x}}}}$ and $\omega\in D_{\bm{\mathsf{x}}}$, $\bar{\bm{\mathsf{x}}}(\omega)={\bm{\mathsf{x}}}(\omega)$. then $\bar{\bm{\mathsf{X}}}={\bm{\mathsf{X}}}$;
    6. (f)
      for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, there is countable ${\bm{\mathsf{X}}}_{0}\subseteq{\bm{\mathsf{X}}}$ such that for all ${\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}$ with ${\bm{\mathsf{x}}}^{\prime}>_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}$ there is ${\bm{\mathsf{x}}}_{0}\in{\bm{\mathsf{X}}}_{0}$ with ${\bm{\mathsf{x}}}^{\prime}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{0}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}$.
@@
The elements of ${\bm{\mathsf{X}}}$ are called random moves. For $\omega\in\Omega$, let $T_{\omega}=\pi^{-1}(\{\omega\})$ and $W_{\omega}$ be the root for $T_{\omega}$. For $A\subseteq\Omega$, let $W_{A}=\bigcup_{\omega\in A}W_{\omega}$ and $F_{A}=\bigcup_{\omega\in A}T_{\omega}$.
Let us discuss this definition in more detail. A stochastic decision forest is a decision forest over a set (1), whose connected components are indexed by $\Omega$ (2), admitting a maximal (3e) and separable (3f) set of sections of moves (3a) called random moves, that form a covering (3b) of $X$, and are compatible with the rooted forest $(F,\supseteq)$ by preserving the order (3c) and identifying roots (3d). As is going to be shown in the sequel, these properties are actually stronger than the definition formally indicates. If one visualises a stochastic decision forest in two dimensions, any tree growing along the vertical “decision path” axis, the forest’s trees being placed along the horizontal “$\Omega$” axis, then random moves are vertical sections partitioning the set of moves in a monotone, decision-theoretically interpretable way. Random moves thus define a structure of “similarity” across different connected components.
Moreover, they can be seen as moves of a rooted decision tree in their own right. Let us make this precise. The first thing to note is that as a consequence of the maximality assumption, random moves cannot be empty:
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be a stochastic decision forest over an exogenous scenario space $(\Omega,\mathscr{A})$. Then, for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, $D_{\bm{\mathsf{x}}}\neq\emptyset$.
Yet, this is not sufficient to ensure that random moves can be distinguished as moves with respect to the order $\geq_{\bm{\mathsf{X}}}$, since we might have to add terminal nodes, if they exist. For example, think of an “infinite centipede” which contains infinitely many maximal chains but only one maximal chain consisting only of moves. Hence, we extend the partial order $\geq_{\bm{\mathsf{X}}}$ to a binary relation $\geq_{\bm{\mathsf{T}}}$ on the disjoint union
${\bm{\mathsf{T}}}={\bm{\mathsf{X}}}\cup\{(\omega,w)\in\Omega\times W\mid\{w\}\in F,\leavevmode\nobreak\ \pi(\{w\})=\omega\}$
by letting, for ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ and $(\omega,w),(\omega^{\prime},w^{\prime})\in{\bm{\mathsf{T}}}\setminus{\bm{\mathsf{X}}}$:
  • 1.
    ${\bm{\mathsf{x}}}\geq_{\bm{\mathsf{T}}}(\omega,w)$ iff $\omega\in D_{\bm{\mathsf{x}}}$ and $w\in{\bm{\mathsf{x}}}(\omega)$;
  • 2.
    $(\omega,w)\geq_{\bm{\mathsf{T}}}(\omega^{\prime},w^{\prime})$ iff $(\omega,w)=(\omega^{\prime},w^{\prime})$;
  • 3.
    $(\omega,w)\ngeq_{\bm{\mathsf{T}}}{\bm{\mathsf{x}}}$.
Note that a pair $(\omega,w)$ as above is nothing else than the set-theoretic function ${\bm{\mathsf{y}}}\colon D_{\bm{\mathsf{y}}}\to F,\omega\mapsto\{w\}$ where $D_{\bm{\mathsf{y}}}=\{\omega\}$. In analogy with random moves and since its image is a terminal node, we call such ${\bm{\mathsf{y}}}$ a random terminal node. The analogy is substantiated by the subsequent proposition which (seemingly) strengthens axiom 3c in the definition above.
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be a stochastic decision forest over some exogenous scenario space $(\Omega,\mathscr{A})$ and ${\bm{\mathsf{T}}}\bullet\Omega=\{({\bm{\mathsf{y}}},\omega)\in{\bm{\mathsf{T}}}\times\Omega\mid\omega\in D_{\bm{\mathsf{y}}}\}$. Then the evaluation map $\operatorname{ev}\colon{\bm{\mathsf{T}}}\bullet\Omega\to F,({\bm{\mathsf{y}}},\omega)\mapsto{\bm{\mathsf{y}}}(\omega)$ is a bijection such that for all $({\bm{\mathsf{y}}}_{1},\omega_{1}),({\bm{\mathsf{y}}}_{2},\omega_{2})\in{\bm{\mathsf{T}}}\bullet\Omega$:
${\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}\text{ and }\omega_{1}=\omega_{2}\quad\Longleftrightarrow\quad{\bm{\mathsf{y}}}_{1}(\omega_{1})\supseteq{\bm{\mathsf{y}}}_{2}(\omega_{2}).$
In order-theoretic terms, the evaluation map defines an order isomorphism between $T\bullet\Omega$, equipped with the order induced by the product of $\geq_{\bm{\mathsf{T}}}$ on ${\bm{\mathsf{T}}}$ and equality on $\Omega$, and $(F,\supseteq)$. Now, as announced, we can make rigorous the sense in that random moves are the moves of a rooted decision tree.

### v2

@@ 1.3 Forests of decision trees
Then $F$ is a decision forest on $V$ iff
@@
    for every $T\in\mathcal{F}$, $(T,\supseteq)$ defines a decision tree on $V_{T}$.
To verify the Propositions 1.4, 1.5, and 1.6, on can cite the results in [3] in combination with Remark 1.2. In view of Theorem 1.7, it suffices, however, to cite the results for the case of rooted trees only. This will be of interest at a later stage when analysing outcome existence and uniqueness for stochastic extensive forms.
@@
Theorem 1.7 offers a way of constructing decision forests from collections of decision trees. Let $\mathcal{F}_{0}$ be a non-empty set of decision trees on sets. For any $T_{0}\in\mathcal{F}_{0}$, let $V_{T_{0}}$ be its root. Then, let
@@
$F=\big{\{}x\times\{T_{0}\}\mid T_{0}\in\mathcal{F}_{0},\,x\in T_{0}\big{\}}.$
Then, $F$ is a decision forest on $V$. The set of connected components is given by
=$\mathcal{F}=\Big{\{}\big{\{}x\times\{T_{0}\}\mid x\in T_{0}\big{\}}\leavevmode\nobreak\ \Big{|}\leavevmode\nobreak\ T_{0}\in\mathcal{F}_{0}\Big{\}}.$
Moreover, Theorem 1.7 states that all decision forests on sets can be represented in this form.
For examples of decision forests on sets, the reader is referred to the following section, where decision forests whose connected components are indexed over the set of exogenous scenarios are considered. [^3] Also note the examples of decision trees on sets in [3] and the more extensive version in [5] in combination with the preceding Remark 1.8.
@@ 2.1 Exogenous scenario spaces
For the remainder of the paper, an exogenous scenario space is a measurable space $(\Omega,\mathscr{E})$ such that $\Omega\neq\emptyset$. [^4] For more information on measure and probability theory, the reader is referred to the introductory texts [11, 32]. $\Omega$ describes the set of possible exogenous scenarios which replace the outcomes possibly generated by the behaviour of a “nature” agent. $\mathscr{E}$ is the set of events alias collections of exogenous scenarios that can be measured by all relevant decision makers. That is, we implictly suppose that decision makers alias agents form beliefs about the probability of events, in the sense of probability theory, and here, $\mathscr{E}$ describes the set of collections of exogenous scenarios that agents assign probabilities to. [^5] These beliefs can be interpreted as representation devices of agents’ preferences over the random outcomes of strategic interaction in terms of expected utility – “random” because of the dependence on exogenous scenarios, which in this particular context play the role of states while outcomes correspond to the consequences, put in Savage’s words, see, e.g. [45, 23, 26].
@@
Why does one suppose $\mathscr{E}$ to be a $\sigma$-algebra? As an algebra of sets, $\mathscr{E}$ can be seen as a Boolean algebra and can therefore model basic logical operations on events. Moreover, in the probabilistic approach, we are only interested in those functions ${\mathbb{P}}\colon\mathscr{E}\to[0,1]$ compatible with the algebra structure [^6] The meaning of the phrase “compatible with the algebra structure” can be reduced to the statement: a) ${\mathbb{P}}(\Omega)=1$, and b) for all disjoint $E_{1},E_{2}\in\mathscr{E}$ we have ${\mathbb{P}}(E_{1}\cup E_{2})={\mathbb{P}}(E_{1})+{\mathbb{P}}(E_{2})$. that are $\sigma$-additive, which can be thought of as a continuity property, and it is a fundamental measure-theoretic result, that, for such ${\mathbb{P}}$, without loss of generality, $\mathscr{E}$ can be supposed to be closed under countable unions, that is, to be a $\sigma$-algebra (see [11, Section 1.5]).
@@
If $\mathscr{E}$ can be generated by a countable partition $\mathscr{P}$ of $\Omega$, any probability measure on it is entirely described through the probabilities on these partitioning events. So, though all elements of $\mathscr{E}$ including unions of partition members are events, one can formally describe the situation by the set $\mathscr{P}$ and a countable family of positive numbers $(p_{E})_{E\in\mathscr{P}}$ adding up to $1$ and indexed by that set. However, in general, $\mathscr{E}$ need not be generated by a partition of $\Omega$. For instance, it may be that all singletons are events and have probability zero, as in the case of the Lebesgue measure on the unit interval $[0,1]$. Hence, from a general perspective, scenarios may be neither a relevant nor a sufficient description of exogenous data for decision makers, but it is the ($\sigma$-)algebra $\mathscr{E}$ of events that these agents are concerned about. Moreover, as famously illustrated by the Lebesgue measure, in case $\Omega$ is uncountable, there is a non-trivial trade-off between the fineness [^7] Recall that a $\sigma$-algebra $\mathscr{E}^{\prime}$ on $\Omega$ is finer than $\mathscr{E}$ iff $\mathscr{E}\subseteq\mathscr{E}^{\prime}$. of $\mathscr{E}$ and the amount of probability measures on it (see [11, Section 1.7]). Hence, to describe the agent’s relation to exogenous data and admissible beliefs about their probability, it is crucial to specify $\mathscr{E}$. This is crucial, not the least because many relevant applications require an uncountable $\Omega$ (with a $\sigma$-algebra $\mathscr{E}$ that cannot be generated by a countable partition), be it the uniform distribution on $[0,1]$, be it continuous martingales omnipresent in finance (see, e.g. [17, 9]).
@@ 2.2 Stochastic decision forests
In light of the preceding argument, we model uncertainty about exogenous events by an exogenous scenario space $(\Omega,\mathscr{E})$ and suppose that some $\omega\in\Omega$ is realised, determining the relevant decision tree without being directly communicated to the decision makers. So, the agents do not necessarily have complete information about exogenous events in $\mathscr{E}$, which tell in which tree they are while making choices. Note that events in $\mathscr{E}$ need not be generated by a partition of $\Omega$. Hence, there must be a structure of similarity among trees that can serve as a consistent basis for exogenous information revelation in an abstract and general probabilistic sense. This distinguishes the present approach from the traditional model of games or decision problems under incomplete information ([27, 28, 29]) or stochastic games ([46], see also [5, Subsection 2.2.2.5]), represented via a randomising nature agent and partition-based information sets with respect to that agent’s past action.
@@
A stochastic decision forest, in short sdf, on an exogenous scenario space $(\Omega,\mathscr{E})$ is a triple $(F,\pi,{\bm{\mathsf{X}}})$ consisting of:
@@
    a decision forest $F$ on some set $W$;
@@
      any element ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ is a section of moves defined on some non-empty event, that is, it is a map ${\bm{\mathsf{x}}}\colon D_{\bm{\mathsf{x}}}\to X$ satisfying $\pi\circ{\bm{\mathsf{x}}}=\operatorname{id}_{D_{\bm{\mathsf{x}}}}$ for some $D_{\bm{\mathsf{x}}}\in\mathscr{E}\setminus\{\emptyset\}$;
    2. (3(b))
      ${\bm{\mathsf{X}}}$ induces a covering of $X$, that is, $\{{\bm{\mathsf{x}}}(\omega)\mid{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}},\,\omega\in D_{\bm{\mathsf{x}}}\}=X$.
@@
The elements of ${\bm{\mathsf{X}}}$ are called random moves. For $\omega\in\Omega$, let $T_{\omega}=\pi^{-1}(\{\omega\})$ and $W_{\omega}$ be the root of $T_{\omega}$. For $E\subseteq\Omega$, let $W_{E}=\bigcup_{\omega\in E}W_{\omega}$ and $F_{E}=\bigcup_{\omega\in E}T_{\omega}$.
In other words, a stochastic decision forest is a decision forest on a set (Axiom 1), whose connected components are indexed by $\Omega$ (Axiom 2), admitting a set of sections of moves (Axiom 33(a)) called random moves that form a covering (Axiom 33(b)) of $X$. As discussed in the remainder of this series of papers, random moves form a flexible basis for a general description of information revelation for different agents, under the innocent hypothesis that there is at least one active agent per move.
If one visualises a stochastic decision forest in two dimensions, any tree growing along the vertical “decision path” axis, the forest’s trees being placed along the horizontal “$\Omega$” axis, then random moves are – roughly speaking – horizontal or diagonal sections covering the set of moves in a decision-theoretically interpretable way. Hence, in a stochastic decision forest, the flow of information can be decomposed into the endogenous movement of agents along the set of random moves caused by decision making and the exogenous information revelation at those random moves. Note also that stochastic decision forests are allowed to vary across scenarios: the trees in $\mathcal{F}$ need not be isomorphic. The crowns may become shallower in some scenarios (because some options are no more available there).
It is possible that random moves as such can be ordered in a way consistent with the ordering on $F$. This means that one can relate these points of exogenous information revelation uniformly by means of the words “before” and “after”. This possibility is necessary for conceiving an agent’s capacity to condition action on knowledge about exogenous information received “earlier”. Note that this order consistency of random moves is not about information sets and, in particular, not to be confounded with the strong order on information sets as introduced in [42]. An example of an order inconsistent stochastic decision forest is presented in the following subsection (namely, Gilboa’s interpretation of the absent-minded driver formulated as a stochastic decision forest). Furthermore, we introduce two non-triviality conditions. In many contexts it seems reasonable to assume first that any root is a move – a situation which (provided sufficient measurability) can always be obtained by eliminating those scenarios $\omega$ with singleton $T_{\omega}$ alias singleton $W_{\omega}$ –, and second that the set of random moves describes the basis of exogenous information most efficiently, in the sense that it cannot be extended without compromising its structure.
Definition 2.10.
Given a stochastic decision forest $(F,\pi,{\bm{\mathsf{X}}})$ on an exogenous scenario space $(\Omega,\mathscr{E})$, let $\geq_{\bm{\mathsf{X}}}$ denote the partial on ${\bm{\mathsf{X}}}$ defined by
${\bm{\mathsf{x}}}_{1}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{2}\quad\Longleftrightarrow\quad\Big{[}D_{{\bm{\mathsf{x}}}_{1}}\supseteq D_{{\bm{\mathsf{x}}}_{2}}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \forall\omega\in D_{{\bm{\mathsf{x}}}_{2}}\colon{\bm{\mathsf{x}}}_{1}(\omega)\supseteq{\bm{\mathsf{x}}}_{2}(\omega)\Big{]}.$
A set $\tilde{\bm{\mathsf{X}}}\subseteq{\bm{\mathsf{X}}}$ of random moves is said order consistent iff for all ${\bm{\mathsf{x}}}_{1},{\bm{\mathsf{x}}}_{2}\in\tilde{\bm{\mathsf{X}}}$:
$\Big{[}\exists\omega\in D_{{\bm{\mathsf{x}}}_{1}}\cap D_{{\bm{\mathsf{x}}}_{2}}\colon\leavevmode\nobreak\ {\bm{\mathsf{x}}}_{1}(\omega)\supseteq{\bm{\mathsf{x}}}_{2}(\omega)\Big{]}\qquad\Longrightarrow\qquad{\bm{\mathsf{x}}}_{1}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{2}.$
A stochastic decision forest $(F,\pi,{\bm{\mathsf{X}}})$
  1. 4(a)
    is said order consistent iff ${\bm{\mathsf{X}}}$ is order consistent;
  2. 4(b)
    is said surely non-trivial iff $(F,\supseteq)$ is everywhere non-trivial;
  3. 4(c)
    that is order consistent, is said maximal iff for every set $\bar{\bm{\mathsf{X}}}$ such that $(F,\pi,\bar{\bm{\mathsf{X}}})$ is an order consistent stochastic decision forest and that is refined by ${\bm{\mathsf{X}}}$ in that for all $\bar{\bm{\mathsf{x}}}\in\bar{\bm{\mathsf{X}}}$ there is $P_{\bar{\bm{\mathsf{x}}}}\subseteq{\bm{\mathsf{X}}}$ with $\bar{\bm{\mathsf{x}}}=\bigcup P_{\bar{\bm{\mathsf{x}}}}$, [^8] According to standard set-theoretic conventions, $\bar{\bm{\mathsf{x}}}=\bigcup P_{\bar{\bm{\mathsf{x}}}}$ means: $\bar{\bm{\mathsf{x}}}$ is a map with domain $\bigcup_{{\bm{\mathsf{x}}}\in P_{\bar{\bm{\mathsf{x}}}}}D_{\bm{\mathsf{x}}}$ and for all ${\bm{\mathsf{x}}}\in P_{\bar{\bm{\mathsf{x}}}}$ and $\omega\in D_{\bm{\mathsf{x}}}$, $\bar{\bm{\mathsf{x}}}(\omega)={\bm{\mathsf{x}}}(\omega)$. we have $\bar{\bm{\mathsf{X}}}={\bm{\mathsf{X}}}$.
As is going to be shown in the sequel, some of these properties are actually stronger than the definition formally indicates. In the case of order consistency, random moves cannot cross each other, and in particular, the root property is compatible with random moves.
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be a stochastic decision forest. Let $D=\{\omega\in\Omega\mid W_{\omega}\in X\}$.
  1. 1.
    If $(F,\pi,{\bm{\mathsf{X}}})$ is order consistent, then, for any ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, ${\bm{\mathsf{x}}}(\omega)$ is a root in $(F,\supseteq)$ for all or no $\omega\in D_{\bm{\mathsf{x}}}$.
  2. 2.
    $D=\Omega$ iff $(F,\pi,{\bm{\mathsf{X}}})$ is surely non-trivial.
  3. 3.
    If $(F,\pi,{\bm{\mathsf{X}}})$ is order consistent and maximal, and $X\neq\emptyset$, then ${\bm{\mathsf{x}}}_{0}\colon D\to X,\,\omega\mapsto W_{\omega}$ is a random move, i.e. ${\bm{\mathsf{x}}}_{0}\in{\bm{\mathsf{X}}}$.
Moreover, in the order consistent, surely non-trivial and maximal case, random moves can be seen as moves of a rooted decision tree in their own right. This makes the endogenous movement of agents along ${\bm{\mathsf{X}}}$ particularly ordered: as said before, it becomes possible to relate two points of exogenous information revelation via “before” and “after”. Let us make this precise. First, we might have to add terminal nodes, if they exist. For example, think of an “infinite centipede” which contains infinitely many maximal chains but only one maximal chain consisting only of moves. Hence, we extend the partial order $\geq_{\bm{\mathsf{X}}}$ to a binary relation $\geq_{\bm{\mathsf{T}}}$ on the disjoint union
${\bm{\mathsf{T}}}={\bm{\mathsf{X}}}\cup\Big{\{}\big{\{}(\omega,\{w\})\big{\}}\leavevmode\nobreak\ \Big{|}\leavevmode\nobreak\ (\omega,w)\in\Omega\times W\colon\{w\}\in F,\,\pi(\{w\})=\omega\Big{\}}$
by letting, for ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ and $(\omega,w),(\omega^{\prime},w^{\prime})\in\Omega\times W$ with $\{w\},\{w^{\prime}\}\in F$ and $\pi(\{w\})=\omega$, $\pi(\{w^{\prime}\})=\omega^{\prime}$:
  • ${\bm{\mathsf{x}}}\geq_{\bm{\mathsf{T}}}\big{\{}(\omega,\{w\})\big{\}}$ iff $\omega\in D_{\bm{\mathsf{x}}}$ and $w\in{\bm{\mathsf{x}}}(\omega)$;
  • $\big{\{}(\omega,\{w\})\big{\}}\geq_{\bm{\mathsf{T}}}\big{\{}(\omega^{\prime},\{w^{\prime}\})\big{\}}$ iff $(\omega,w)=(\omega^{\prime},w^{\prime})$;
  • $\big{\{}(\omega,\{w\})\big{\}}\ngeq_{\bm{\mathsf{T}}}{\bm{\mathsf{x}}}$.
Note that a set $\big{\{}(\omega,\{w\})\big{\}}$ as above is nothing else than the set-theoretic function ${\bm{\mathsf{y}}}\colon D_{\bm{\mathsf{y}}}\to F$ with $D_{\bm{\mathsf{y}}}=\{\omega\}$ and ${\bm{\mathsf{y}}}(\omega)=\{w\}$. In analogy with random moves and since its image is a terminal node, we call such ${\bm{\mathsf{y}}}$ a random terminal node. The analogy is substantiated by the subsequent proposition which (seemingly) strengthens Axiom 4(a) in Definition 2.10 above.
@@
Let $(F,\pi,{\bm{\mathsf{X}}})$ be an order consistent stochastic decision forest on some exogenous scenario space $(\Omega,\mathscr{E})$ and ${\bm{\mathsf{T}}}\bullet\Omega=\{({\bm{\mathsf{y}}},\omega)\in{\bm{\mathsf{T}}}\times\Omega\mid\omega\in D_{\bm{\mathsf{y}}}\}$. Then the evaluation map $\operatorname{ev}\colon{\bm{\mathsf{T}}}\bullet\Omega\to F,({\bm{\mathsf{y}}},\omega)\mapsto{\bm{\mathsf{y}}}(\omega)$ is a bijection such that for all $({\bm{\mathsf{y}}}_{1},\omega_{1}),({\bm{\mathsf{y}}}_{2},\omega_{2})\in{\bm{\mathsf{T}}}\bullet\Omega$:
$\Big{[}{\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}\text{ and }\omega_{1}=\omega_{2}\Big{]}\quad\Longleftrightarrow\quad{\bm{\mathsf{y}}}_{1}(\omega_{1})\supseteq{\bm{\mathsf{y}}}_{2}(\omega_{2}).$
In order-theoretic terms, the evaluation map defines an order isomorphism between ${\bm{\mathsf{T}}}\bullet\Omega$, equipped with the order induced by the product of $\geq_{\bm{\mathsf{T}}}$ on ${\bm{\mathsf{T}}}$ and equality on $\Omega$, and $(F,\supseteq)$. Now, as announced, we can make rigorous the sense in that random moves are the moves of a rooted decision tree, under the assumptions presented in Definition 2.10.